OpenStudy (mendicant_bias):

I'm trying to figure out how to reindex a series into another one. (Math posted in a moment.)

OpenStudy (mendicant_bias):

Original series: \[\sum_{n = 1}^{\infty} (\frac{ n }{ 2n+3 })^{n}\]

OpenStudy (mendicant_bias):

I need to reindex that into...\[\sum_{n = ?}^{\infty}(\frac{ n }{ 2n })^{n}\]

OpenStudy (mendicant_bias):

Most of the problems that I have seen involving reindexing have reduced the index, but it looks like I'll need to do the reverse here, and I'm not so sure how to do that satisfactorily.

OpenStudy (zzr0ck3r):

i dont think you can, not by just changing n

OpenStudy (anonymous):

You just take a rubix cube and solve it.